# Convex Approximation of Probabilistic Reachable Sets from Small Samples Using Self-supervised Neural Networks

### v1

@@ III-B Expert method
The expert method is proposed in [40]. The expert method can generate the PRS of an uncertain dynamic system. This data-driven method builds on a kernel density estimator (KDE) that can generate the PRS, accelerated by a fast Fourier transform (FFT). These KDE values are then utilized within an Integer Linear Programming (ILP) framework to find a minimal parallelogram box that approximates the probabilistic geofence. This ILP solution effectively encapsulates the area where the state is expected to stay with high confidence.
@@
The ILP Heuristic Algorithm approximates solutions for integer linear programming problems using heuristics. It starts with a weight matrix $\omega_{ij}$ and a zero matrix $z_{ij}$, setting initial indices based on $\omega_{ij}$’s maximum values. The algorithm loops to update $z_{ij}$ until the product of $z_{ij}$ and $\omega_{ij}$ meets a threshold $\alpha$. Each iteration zeros out the maximum $\omega_{ij}$ and adjusts boundary indices. The corresponding $z_{ij}$ entry is set to one within these boundaries, yielding the matrix as an efficient approximation of the ILP solution. The expert convex generated is shown in Fig. 2. The black points represent the distribution and purple convex covers a desired percent of the distribution.
III-C Training dataset generation
Theoretically, there are unlimited correct convex approximations for PRS given a probability distribution. At the same time, the expert method introduced in the previous section can generate a very accurate convex approximation if enough samples are given. Therefore, for each distribution $P$ and each disired confidence level $\alpha$, we can generate a set of correct convex approximation $\mathbf{\tilde{\mathbf{S}}}_{p,\alpha}=\{\tilde{\mathcal{S}}_{P,\alpha}^{i}\}_{i=1}^{n}$. In this paper, the correct convex approximation has 1% error toleration. Therefore, we consider the approximation is correct if $\alpha+0.01\geq\operatorname{Pr}\left(\boldsymbol{s}\in\tilde{\mathcal{S}}_{\alpha}\right)\geq\alpha$. Each correct convex approximation $\tilde{\mathcal{S}}_{P,\alpha}$ is defined by four vertices $\tilde{\mathcal{V}}_{P,\alpha}$. Then the set of vertices of the correct convex approximation is denoted as $\mathbf{\tilde{\mathbf{V}}}_{p,\alpha}=\{\tilde{\mathcal{V}}_{P,\alpha}^{i}\}_{i=1}^{n}$. The usage of the generated correct convex approximation will be introduced in the method section below.
IV Method
IV-A Self-supervised learning
Figure 3: Masked distribution: we only show a few samples to the neural networks, and ask neural networks to generate the corresponding PRS.
Previous research [33] has proven that the transformer can process the visible (unmasked) parts of the data, reconstructing the entire input from the encoded, incomplete data (masked data). Therefore, if we only have a few samples from the distribution, the transformer may be able to reconstruct the entire distribution.
As the previous section III, we have the vertices of expert convex approximation $\mathbf{\tilde{\mathbf{V}}}_{p,\alpha}$. Fig.3 shows the examples of how we mask the distribution. During the training, we randomly mask most of the sample points, only leaving a few samples visible to the transformer, and ask the transformer to generate the corresponding PRS.
With this self-supervised learning strategy, we are able to show many sets of visible samples of one distribution to the neural networks. When we show the stranger samples to the trained neural networks, the trained neural network is able to identify the distribution by recognizing and relating the visible samples to the samples the neural networks have seen during the training process.
IV-B Multi-label learning and loss function
Figure 4: Example of multi-label learning
As introduced in section II, one distribution can have multiple correct convex approximations. Therefore, we can use multi-label learning instead of single-label learning. Multi-label learning not only reduces computational overhead but also allows the model to share learned features across multiple labels, potentially leading to better generalization. Fig. 4 shows an example of multi-label learning. The neural network generated an output based on the current weights of the neural network. Then the losses between each expert and the output are calculated and the smallest loss will be used to train the neural network. For the real training, we will have more expert PRS samples. Like many regression machine learning research studies, we use the MSE loss function as shown in Eq. 7. MSE loss can help the neural network minimize the average error between the expert and the output.
@@ IV-B Multi-label learning and loss function
$\text{Loss}=\min_{j\in[1,10]}{\frac{1}{8}\sum_{i=1}^{8}(\hat{y}_{ij}-y_{i})^{2}}\quad\text{where}\quad\hat{y}_{ij}\in\mathbf{\tilde{\mathbf{V}}}_{p,\alpha}$ (7)
@@
IV-C Neural networks architecture
The overall architecture of the neural network we used is shown in Fig. 5, which is designed based on our previous work [43]. The input to the neural network consists of sampled points and the desired confidence level. Before sending these inputs to the neural network, we normalize them first. The confidence level and the sample points are then embedded in a higher dimension using a linear layer. The embedded items are concatenated together. These concatenated embeddings are then sent to the transformer layer. In the transformer, the channel representing the confidence level interacts with all the point channels. Finally, we send the confidence level channel to an output layer, resulting in four vertices of the convex hull.
@@ IV-C Neural networks architecture
IV-D Input normalization, embedding, and concatenate
Because sample points are drawn from different distributions, they have very different means and variances, which can cause challenges for the training process of neural networks. Normalizing those points before feeding them into the neural network is very crucial for several reasons. First, if certain inputs have larger values and higher variance than others, they may dominate the training process. For example, if one target (or feature) has a very large value compared to others, its gradient will be disproportionately large. Second, normalization can accelerate the training process. When the scales of the inputs are not consistent, the model might spend more time adjusting weights to accommodate the varying scales of input data rather than learning meaningful patterns. Third, normalization can prevent overfitting by reducing bias toward dominant features and indirect regularization effects.
@@ IV-E Transformer
For this work, we only use the transformer encoder [16] since we output the vertex at once instead of one by one. The encoder is composed of a stack of 6 identical layers, each layer has two sub-layers, a multi-head self-attention mechanism, and the feedforward network. We also kept the residual connection and layer normalization. Compared to other popular layers, such as the recurrent and convolutional layers, self-attention has three main advantages. First, self-attention has lower computational complexity per layer. Second, it requires fewer sequential operations since most computations can be parallelized. The third advantage is that self-attention can learn the dependencies between two points that are far from each other. For example, if we use CNNs, we may input points patch by patch, then it is difficult to learn the relationship between the first patch and the last patch. In contrast, each point interacts with all the other points at the same time in self-attention.
IV-F Output
In the previous masked autoencoder research. Mask tokens are introduced after the encoder, so the decoder will know which missed position is required to be generated. However, unlike reconstructing assignments, we do not have mask tokens. We do not know the visible samples are from which part. In fact, unlike CV and NLP problems, the sequence does not matter in our problem. For example, ”The dog chased the cat” and ”The cat chased the dog” are two completely different sentences, even though they have exactly the same words. In contrast, in our problem, the point set [(1, 1), (2, 2)], and [(2,2), (1, 1)] have the same distribution. Therefore, no additional global context is required. Furthermore, the goal of the proposed neural network is to generate a convex bound instead of reconstructing the whole distribution. All in all, a simple output layer is enough to generate the convex bound with the features captured from the visible samples.
@@ IV-F Output
We use a multilayered output like ViT [15] instead of a single linear layer like the original transformer. The multi-layer output can further process the features extracted from the points channels and refine the final classification. We only used the transformed confidence level channel to output. The transformer allows the confidence level channel to attend to all other channels (embedded points), gathering contextual information from all the points. Finally, we need to denormalize the output, since the points we want to bound are normalized.
IV-G Parameters
@@
For the transformer, embedding dimension $N_{emd}$ is 864. We use 6 attention heads, so each head processes 144 embedding dimensions of each channel. The number of attention layer is also 6. Optimization is performed using the Adam optimizer with a dynamic learning rate to ensure gradual convergence. Additionally, a dropout rate of 0.2 is applied to prevent overfitting by randomly omitting certain neurons during training.
V Result
(a) 2d-binominal distribution
(b) Fanshape distribution
(c) GMR trajectory predictor distribution
@@ V Result
Figure 7: Example outputs if three methods take the same visible points, visible points are enlarged as yellow points for greater visibility.
V-A Experiment setting
We compare the proposed method with the ILP Heuristic (ILP), which is the expert method, and the bounding box (BB) method [45][46], a popular approach to finding a convex approximation of the PRS. Both baseline methods are KDE-based. Fig. 7 are examples of point distribution and convex hulls generated by different methods given the same visible points. All the experiences are run in Intel(R) Core(TM) i9-12900KF and NVIDIA GeForce RTX 3090.
V-A1 Training dataset
In each training iteration, we randomly assign ten sets of expert vertices as the target and randomly pick 100 samples (300 for the GMR trajectory predictor) from the desired distribution as the neural network’s input. To accelerate the training process, we prepare 100 sets of expert vertices for each desired confidence level and prepare 320,000 sets of samples before training begins. To make the required expert vertices, it will cost experts about 24 hours for all distributions. In addition to expert vertices generating, we also need to generate training input samples before training. The sample generation time varies across different distributions. For the binomial distribution, it takes only about 2 minutes to generate 320,000 sets of samples. For the fanshape distribution, the process takes approximately 40 minutes to produce the same number of sample sets. In the case of the GMR trajectory predictor, it requires around 30 hours to generate 320,000 sets of samples. Generally, we can prepare the training dataset for a known distribution within a few days. It supports that the Advantage 4 of the proposed method, the training dataset is easy to make. Additionally, the proposed method exhibits Advantage 3: all PRS are guaranteed to be convex and linear, as the neural network is specifically designed to generate only the vertices of convex sets.
V-A2 PRS generating
We tested the trained neural networks by presenting them with 5,000 random tasks (1000 for each $\alpha$) in each distribution. Each task contains 100 (300 for GMR) points of the distribution and has never been seen by the neural network. The neural network will generate the corresponding PRS for those points. At the same time, we also generate PRS with BB and ILP given the same points as the baseline. However, since BB and ILP are not learning-based and do not require a training process, comparing them directly may not be entirely fair. To adjust for this, we also compared the results using PRS generated by BB and ILP with 20 times more samples to thoroughly evaluate our proposed method. Besides the baseline method, we also run the MINLP Optimal Algorithm [47] to determine the smallest area of a convex satisfies the coverage requirement.
V-A3 PRS evaluating
To evaluate a PRS, we first need to determine the coverage. To represent the real distribution, we sampled 100,000 points from the desired distribution. The resulting coverage was determined by counting the percentage of these sampled points that fell within the generated PRS.
Mean Squared Error(MSE) is the metric we used to evaluate the accuracy with which the PRS approximates the desired confidence level:
@@ V-A3 PRS evaluating
$\alpha$ is the coverage of the generated PRS and $\hat{\alpha}_{i}$ is the desired coverage. n is 1000 because we test 1000 times for each $\alpha$. We also recorded the minimum and maximum coverage values to evaluate performance stability. While a low MSE indicates that the PRS is accurate, it does not guarantee correctness. Therefore, it is essential to compare the percentage of instances where the PRS exceeds the desired coverage. We called it the correctness rate.
V-A4 Generating time
Generating time is a crucial property in deciding whether a method can be used in online applications. Therefore, we compare the time to generate one correct PRS by each method.
V-B Compare coverage accuracy, correctness, and area to baseline
2d-binomial NN(our method) BB(100) ILP(100)
confidence level MSE(range) MSE(range) MSE(range)
@@ V-B Compare coverage accuracy, correctness, and area to baseline
TABLE I: MSE and range on 2d-binomial distribution distribution compare to baseline with a small number of sample
fanshape NN(our method) BB(100) ILP(100)
confidence level MSE(range) MSE(range) MSE(range)
99% 0.98%(99.96% - 99.99%) 1.33%(98.03% - 100%) 1.12%(97.86% - 100%)
TABLE II: MSE and range on fanshape distribution distribution compare to baseline with a small number of sample
GMR trajectory predictor NN(our method) BB(300) ILP(300)
confidence level MSE(range) MSE(range) MSE(range)
TABLE III: MSE and range on GMR trajectory predictor distribution compare to baseline with a small number of sample
2d-binomial NN(our method) BB(2000) ILP(2000)
confidence level MSE(range) MSE(range) MSE(range)
@@
TABLE IV: MSE and range on 2d-binomial distribution distribution compare to baseline with a large number of sample
fanshape NN(our method) BB(2000) ILP(2000)
confidence level MSE(range) MSE(range) MSE(range)
99% 0.98%(99.96% - 99.99%) 0.99%(99.96% - 100%) 0.98%(99.19% - 100%)
TABLE V: MSE and range on fanshape distribution distribution compare to baseline with a large number of sample
GMR trajectory predictor NN(our method) BB(6000) ILP(6000)
confidence level MSE(range) MSE(range) MSE(range)
TABLE VI: MSE and range on GMR trajectory predictor distribution compare to baseline with a large number of sample
2d-binomial correctness rate
confidence level NN BB(100) BB(2000) ILP(100) ILP(2000) optimal(10000)
@@
confidence level PRS area
@@
TABLE VII: correctness rate/PRS area for each desired confidence level on 2d-binomial distribution
fanshape correctness rate
confidence level NN BB(100) BB(2000) ILP(100) ILP(2000) optimal(10000)
80% 98.7% 100% 100% 57.7% 79.8% 100%
85% 99.8% 100% 100% 52.4% 64.4% 100%
90% 99.7% 99.8% 100% 52.8% 82.3% 100%
95% 100% 99.4% 100% 74.8% 98.9% 100%
99% 100% 61% 100% 59.7% 100% 100%
confidence level PRS area
80% 1415.07 2342.52 2309.52 1783.52 1514.04 1172.71
85% 1762.30 2790.63 2815.55 2138.71 1902.27 1579.52
90% 2229.01 3367.96 3470.37 2652.60 2365.27 2107.79
95% 3530.87 4525.46 4576.97 3382.33 3374.61 2742.49
99% 4833.31 5749.98 6127.58 4967.29 4820.67 3711.22
TABLE VIII: correctness rate/PRS area for each desired confidence level on fanshape distribution
GMR trajectory predictor correctness rate
confidence level NN BB(300) BB(6000) ILP(300) ILP(6000) optimal(30000)
80% 93.1% 100% 100% 10.8% 3.9% 100%
85% 95.2% 100% 100% 27.9% 16.2% 100%
90% 99.8% 100% 100% 69.2% 55.8% 100%
95% 95.6% 100% 100% 68.2% 65.4% 100%
99% 100% 99.7% 100% 56.5% 75.8% 100%
confidence level PRS area
80% 2.75E-3 3.64E-3 3.46E-3 2.59E-3 2.39E-3 2.15E-3
85% 3.44E-3 4.41E-3 4.35E-3 2.99E3 2.89E-3 2.39E-3
90% 4.41E-3 7.37E-3 5.81E-3 3.83E-3 3.67E-3 2.99E-3
95% 5.69E-3 7.89E-3 7.98E-3 5.03E-3 4.87E-3 3.95E-3
99% 9.21E-3 11.19E-3 11.16E-3 8.01E-3 7.63E-3 6.46E-3
TABLE IX: correctness rate/PRS area for each desired confidence level on GMR trajectory predictor distribution
First, we aim to evaluate whether the proposed method can effectively produce accurate PRS. Table III presents the MSE and range comparisons between the proposed method and two baseline approaches, given the exact same visible points on the 2d binomial distribution. Our method consistently achieves lower MSE and a narrower range than the baseline methods, except at a 99% confidence level, where its performance closely matches that of ILP. Table III shows the MSE and range results on the fanshape distribution. The proposed method demonstrates exceptional performance, with significantly lower MSE and a tighter range across all confidence levels. Finally, Table III corroborates the findings from the previous tables, further showcasing that the proposed method achieves substantially lower MSE and a consistently tighter range, underscoring its superior performance. All in all, we found out that our method can produce very accurate PRS. All MSE is lower than 1%, with smaller intervals, indicating our method is very stable and consistent. In contrast, KDE-based baseline methods have higher MSE and variance especially when the desired confidence level is lower. On the other hand, it is not entirely fair to compare KDE-based methods directly, as they do not require a training process. Theoretically, estimating a distribution accurately with a small number of samples is nearly impossible. Therefore, it is also necessary to evaluate whether the trained neural networks can outperform the baseline methods when a sufficient number of samples are provided. Tabel VI VI VI present the MSE and range comparisons between the proposed method (can have only 100 visible samples) and two baseline methods, given 2000 visible samples on three test distributions. While the results of the two baseline methods show slight improvements, they still fall significantly short of the performance achieved by our method.
Second, we want to ensure the convex bound generated by the neural networks is correct and useful. Table IX shows the performance of the neural network on 2d-binomial distribution. The neural network gets at least 98.7% correctness rate and only misses a few tests. Meanwhile, the average area we got is not significantly larger than the optimal area especially when $\alpha$ is 80%. As $\alpha$ increases, the difference between our area and the optimal area also increases. To guarantee reaching the confidence level, the neural networks tend to cover slightly more of the distribution. Since the distribution samples become sparser in the area far from the mean, even a small increase in coverage results in a significant increase in the area. Compared to the bounding box, while it may achieve a higher correctness rate due to its highly conservative nature, its area is significantly larger than that of the proposed method. Meanwhile, although the area of ILP is similar to that of NN, its correctness rate is significantly lower. Table IX shows the performance of the neural network on fanshape distribution. The neural network achieves a near-perfect correctness rate of close to 100%. Additionally, the area comparison follows a similar trend to that observed in the 2d binomial distribution. Table. IX shows the performance of the neural network on GMR trajectory predictor distribution. Although the correctness rate is not as high as in the previous tables, it still reaches a minimum of 93.1%. Meanwhile, the area remains superior to the Bounding Box, comparable to ILP, and only slightly larger than the optimal solution.
In summary, the proposed neural network demonstrates consistently high correctness rates across various distributions, ranging from at least 93.1% to nearly 100%, while maintaining a compact area that is only marginally larger than the optimal solution. This balance between correctness and efficiency underscores the utility of the neural network-generated convex sets. The result supports that the Advantage 1 of the proposed method, NN can generate useful PRS with only a few samples.
V-C Generating speed

### v2

@@ III-A Supervised learning and expert method
In this paper, we employ the expert method, as proposed in our previous work [26]. The expert method can generate the probabilistic obstacle space of an uncertain object. This data-driven method builds on a kernel density estimator (KDE) that can generate the probabilistic obstacle space, accelerated by a fast Fourier transform (FFT). These KDE values are then utilized within an Integer Linear Programming (ILP) framework to find a minimal parallelogram box that approximates the probabilistic obstacle space. This ILP solution effectively encapsulates the area where the state is expected to stay with high confidence.
@@
The ILP Heuristic Algorithm approximates solutions for integer linear programming problems using heuristics. It starts with a weight matrix $\omega_{ij}$ and a zero matrix $z_{ij}$, setting initial indices based on $\omega_{ij}$’s maximum values. The algorithm loops to update $z_{ij}$ until the product of $z_{ij}$ and $\omega_{ij}$ meets a threshold $\alpha$. Each iteration zeros out the maximum $\omega_{ij}$ and adjusts boundary indices. The corresponding $z_{ij}$ entry is set to one within these boundaries, yielding the matrix as an efficient approximation of the ILP solution.
After obtaining the expert polygon vertices $\tilde{\mathcal{V}}$, we prepare the input data for network training. Fig. 4 illustrates how different subsets of samples are presented to the neural network during training. At each training step, a small random subset of samples is selected, and the transformer is trained to generate the corresponding polygon vertices. Through this supervised learning process, the network is repeatedly exposed to multiple partial views of the same obstacle, learning to associate incomplete sample sets with the full probabilistic obstacle space. At inference time, when only a limited number of samples are available, the trained network can infer the underlying distribution of the obstacle space by leveraging the patterns it has learned from these partial-to-full mappings.
III-B Multi-label learning and loss function
Figure 5: Multi-label learning: during training, the network is guided to produce outputs that align with the closest expert solution.
Recall the non-uniqueness of confidence regions [27]: for any given distribution, there exist infinitely many subsets that capture the same probability mass. Consequently, for each uncertain obstacle, there are infinitely many valid probabilistic obstacle spaces. This motivates the use of multi-label learning rather than single-label learning. Multi-label learning not only reduces computational overhead but also allows the model to share learned features across multiple labels, potentially leading to better generalization. Fig. 5 illustrates this process. Given the current network parameters, the neural network produces an output polygon. The loss is then computed between this output and each available expert probabilistic obstacle space. The smallest loss among them is used to update the network weights. In practice, multiple expert probabilistic obstacle spaces are provided for each obstacle during training. Similar to many regression-based approaches, we employ the mean squared error (MSE) loss, as shown in Eq. 5, which minimizes the average discrepancy between the network output and the expert labels.
@@ III-B Multi-label learning and loss function
$\text{Loss}=\min_{j\in[1,10]}{\frac{1}{8}\sum_{i=1}^{8}(\hat{y}_{ij}-y_{i})^{2}}\quad\text{where}\quad\hat{y}_{ij}\in\mathbf{\tilde{\mathcal{V}}}$ (5)
@@
III-C Neural networks architecture
The overall architecture of the neural network we used is shown in Fig. 6, which is designed based on our previous work [28]. The input to the neural network consists of sampled points and the desired confidence level. Before sending these inputs to the neural network, we normalize them first. The confidence level and the sample points are then embedded in a higher dimension using a linear layer. The embedded items are concatenated together. These concatenated embeddings are then sent to the transformer layer. In the transformer, the channel representing the confidence level interacts with all the point channels. Finally, we send the confidence level channel to an output layer, resulting in four vertices of the convex polygon.
@@ III-C Neural networks architecture
III-D Input normalization, embedding, and concatenate
Because sample points are drawn from different tasks, they have very different means and variances, which can cause challenges for the training process of neural networks. Normalizing those points before feeding them into the neural network is very crucial for several reasons. First, if some inputs have larger values and higher variance than others, they may dominate the training process. For example, if one target (or feature) has a very large value compared to others, its gradient will be disproportionately large. Second, normalization can accelerate the training process. When the scales of the inputs are not consistent, the model might spend more time adjusting weights to accommodate the varying scales of input data rather than learning meaningful patterns. Third, normalization can prevent overfitting by reducing bias toward dominant features and indirect regularization effects.
@@ III-E Transformer
For this work, we only use the transformer encoder [30] because our task requires predicting all polygon vertices at once, instead of generating them sequentially. The encoder is composed of a stack of 6 identical layers, each layer has two sub-layers, a multi-head self-attention mechanism, and a feedforward network. We also kept the residual connection and layer normalization. Compared to other popular layers, such as the recurrent and convolutional layers, self-attention has three main advantages. First, self-attention has lower computational complexity per layer. Second, it requires fewer sequential operations since most computations can be parallelized. The third advantage is that self-attention can learn the dependencies between two points that are far from each other. For example, if we use CNNs, we may input points patch by patch, then it is difficult to learn the relationship between the first patch and the last patch. In contrast, each point interacts with all the other points at the same time in self-attention.
III-F Output
In prior masked autoencoder research, mask tokens are introduced after the encoder so that the decoder can identify which missing positions need to be reconstructed. Our setting is fundamentally different. Unlike reconstruction tasks, we do not use mask tokens, since the visible samples do not carry positional information about which part of the obstacle space they come from. Moreover, sequence order is irrelevant in our problem. In natural language processing (NLP), for example, the sentences ‘The dog chased the cat’ and ‘The cat chased the dog’ convey completely different meanings despite containing the same words. In contrast, for our problem, the point sets (1,1),(2,2){(1,1),(2,2)} and (2,2),(1,1){(2,2),(1,1)} represent the same distribution. Consequently, no additional global context is required. Furthermore, the objective of our neural network is not to reconstruct the full underlying distribution but to generate a convex bounding region. Thus, a simple output layer suffices to produce the convex bound from the features extracted by the encoder.
@@ III-F Output
We use a multilayered output like ViT [31] instead of a single linear layer like the original transformer. The multi-layer output can further process the features extracted from the point channels and refine the final classification. We only used the transformed confidence level channel for output. The transformer allows the confidence level channel to attend to all other channels (embedded points), gathering contextual information from all the points. Finally, we need to denormalize the output, since the points we want to bound are normalized.
III-G Parameters
@@
For the transformer, the embedding dimension $N_{emd}$ is 864. We use 6 attention heads, so each head processes 144 embedding dimensions of each channel. The number of attention layer is also 6. Optimization is performed using the Adam optimizer with a dynamic learning rate to ensure gradual convergence. Additionally, a dropout rate of 0.2 is applied to prevent overfitting by randomly omitting certain neurons during training.
IV Result
(a) motion uncertainty
(b) behavior uncertainty
(c) sensor uncertainty
@@ IV Result
Figure 8: Black points are sampled from the real obstacle distribution. Colorful polygons are example outputs when the three methods take the same yellow visible points. Yellow points are enlarged for visibility.
IV-A Experiment setting
To evaluate the effectiveness of the proposed method, we conduct experiments on three types of uncertain obstacles: flying obstacles subject to motion uncertainty, obstacles with uncertain destinations, and obstacles whose positions are estimated using noisy sensor measurements. We feed the neural network with samples drawn from the true obstacle distribution and desired confidence level $\alpha$, then evaluate whether the generated probabilistic obstacle space achieves the desired coverage, i.e., whether it bounds the specified proportion of the distribution. Obstacles of the same type share the same underlying distribution but may exhibit different geometric shapes.
IV-A1 Training
We trained a neural network that can work with varying values of $\alpha$, where $\alpha$ can be 80%, 85%, 90%, 95% and 99%. To accelerate the training process, we prepare 100 sets of expert vertices for each desired confidence level and prepare 320,000 sets of 100 samples before training began. In each training iteration, we randomly select one set of samples and one of $\alpha$ as the neural network’s input, then randomly assign ten sets of corresponding expert vertices as the target. We can prepare the training dataset for a known uncertain obstacle within a few hours. It supports the Advantage 3 of the proposed method, the training dataset is easy to make.
IV-A2 Interface
We tested the trained neural networks by presenting them with 5,000 random tasks (1000 for each $\alpha$) from each type of uncertain obstacle. Each task contains 100 samples randomly generated from the real distribution of the uncertain obstacle and has never been seen by the neural network. The neural network generates the corresponding $o^{\alpha}$ for those samples.
IV-A3 evaluating
To evaluate the $o^{\alpha}$ generated by the neural network, we first need to determine the coverage. We approximate the true obstacle distribution using 100,000 samples, as the underlying distributions are typically empirical. The resulting coverage was computed as the percentage of samples that fell within the generated $o^{\alpha}$. We use mean squared error (MSE) to quantify the accuracy of the $o^{\alpha}$:
@@ IV-A2 Interface
$\alpha_{i}$ is the real coverage of the generated $o^{\alpha}$ and $\hat{\alpha}_{i}$ is the desired confidence level (CL). n is 1000 because we test 1000 times for each $\alpha$. We also recorded the minimum and maximum coverage values to evaluate performance stability. While a low MSE indicates that the probabilistic obstacle space is accurate, it does not guarantee correctness. Therefore, it is essential to compare the percentage of instances where $\alpha_{i}$ exceeds $\hat{\alpha}_{i}$. We called it the correctness rate:
$\text{Correctness Rate}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\!\left[\alpha_{i}>\hat{\alpha}_{i}\right],$ (8)
where $1[\bullet]$ is the indicator function(1 if the condition is true, 0 otherwise).
IV-A4 baseline
We compare the proposed method with the ILP Heuristic (ILP), which is the expert method, and the bounding box (BB) method [32][33], a popular approach to finding a polygonal approximation of the obstacle space. Both baseline methods are KDE-based. We generate $o^{\alpha}$ with BB and ILP for the same obstacle as the baseline. Fig. 8 are examples of obstacle distribution and $o^{\alpha}$ generated by different methods given the same samples. However, since BB and ILP are not learning-based methods and do not require a training process, a direct comparison may not be entirely fair. To address this, we also provide the baseline methods with 20 times more samples to enable a more thorough evaluation of our approach. In addition to these baselines, we implement the MINLP Optimal Algorithm [34] to compute the smallest convex region that satisfies the coverage requirement. All the experiences are run in Intel(R) Core(TM) i9-12900KF and NVIDIA GeForce RTX 3090.
TABLE I: MSE and range on motion uncertainty distribution compared to baseline with a small number of samples
NN(our method) BB(100) ILP(100)
CL MSE(range) MSE(range) MSE(range)
99% 0.98%(99.96% - 99.99%) 1.33%(98.03% - 100%) 1.12%(97.86% - 100%)
TABLE II: MSE and range on motion uncertainty compare to baseline with a large number of sample
NN(our method) BB(2000) ILP(2000)
CL MSE(range) MSE(range) MSE(range)
99% 0.98%(99.96% - 99.99%) 0.99%(99.96% - 100%) 0.98%(99.19% - 100%)
TABLE III: correctness rate/Probabilistic obstacle space area for each desired confidence level on motion uncertainty
Correctness rate
CL NN BB(100) BB(2000) ILP(100) ILP(2000) optimal(10000)
80% 98.7% 100% 100% 57.7% 79.8% 100%
85% 99.8% 100% 100% 52.4% 64.4% 100%
90% 99.7% 99.8% 100% 52.8% 82.3% 100%
95% 100% 99.4% 100% 74.8% 98.9% 100%
99% 100% 61% 100% 59.7% 100% 100%
CL Probabilistic obstacle space area
80% 1415.07 2342.52 2309.52 1783.52 1514.04 1172.71
85% 1762.30 2790.63 2815.55 2138.71 1902.27 1579.52
90% 2229.01 3367.96 3470.37 2652.60 2365.27 2107.79
95% 3530.87 4525.46 4576.97 3382.33 3374.61 2742.49
99% 4833.31 5749.98 6127.58 4967.29 4820.67 3711.22
TABLE IV: MSE and range on behavior uncertainty compare to baseline with a small number of sample
NN(our method) BB(300) ILP(300)
CL MSE(range) MSE(range) MSE(range)
TABLE V: MSE and range on behavior uncertainty compare to baseline with a large number of sample
NN(our method) BB(6000) ILP(6000)
CL MSE(range) MSE(range) MSE(range)
TABLE VI: correctness rate/Probabilistic obstacle space area for each desired confidence level on behavior uncertainty
Correctness rate
CL NN BB(300) BB(6000) ILP(300) ILP(6000) optimal(30000)
80% 93.1% 100% 100% 10.8% 3.9% 100%
85% 95.2% 100% 100% 27.9% 16.2% 100%
90% 99.8% 100% 100% 69.2% 55.8% 100%
95% 95.6% 100% 100% 68.2% 65.4% 100%
99% 100% 99.7% 100% 56.5% 75.8% 100%
CL Probabilistic obstacle space area
80% 2.75E-3 3.64E-3 3.46E-3 2.59E-3 2.39E-3 2.15E-3
85% 3.44E-3 4.41E-3 4.35E-3 2.99E3 2.89E-3 2.39E-3
90% 4.41E-3 7.37E-3 5.81E-3 3.83E-3 3.67E-3 2.99E-3
95% 5.69E-3 7.89E-3 7.98E-3 5.03E-3 4.87E-3 3.95E-3
99% 9.21E-3 11.19E-3 11.16E-3 8.01E-3 7.63E-3 6.46E-3
TABLE VII: MSE and range on sensor uncertainty distribution compare to baseline with a small number of sample
NN(our method) BB(100) ILP(100)
CL MSE(range) MSE(range) MSE(range)
@@
TABLE VIII: MSE and range on sensor uncertainty distribution compare to baseline with a large number of sample
NN(our method) BB(2000) ILP(2000)
CL MSE(range) MSE(range) MSE(range)
@@
TABLE IX: correctness rate/Probabilistic obstacle space area for each desired confidence level on sensor uncertainty
Correctness rate
CL NN BB(100) BB(2000) ILP(100) ILP(2000) optimal(10000)
@@
CL Probabilistic obstacle space area
@@
IV-B Motion Uncertainty
For motion uncertainty, we evaluate flying obstacles subject to stochastic deviations in their trajectories. We simulate a flight using the Dubins vehicle model, where both speed and heading angle are sampled from a truncated Gaussian distribution. The resulting obstacle distribution for this simulated flight is shown in Fig. 8a.
[…]
Correctness rates in Table III further highlight the robustness of our approach. NN achieves above 98% correctness at all CLs, while ILP often falls below 60% with limited samples. BB can achieve high correctness, but only at the cost of significantly larger probabilistic obstacle space areas, which reduces path planning efficiency. In contrast, NN produces obstacle spaces closer to the optimal size, especially at higher CLs.
IV-C Behavior Uncertainty
For behavior uncertainty, we evaluate flying obstacles with uncertain destinations. The real obstacle distribution is generated using a Gaussian Mixture Regression (GMR) trajectory predictor, which estimates the probability distribution of a flight’s future positions in terminal airspace [35]. Given the past trajectory, the predictor models multiple possible future routes as a mixture of Gaussians. The resulting obstacle distribution for this predicted flight is shown in Fig. 8b.
Results under behavior uncertainty (Tables VIVI) reveal the strength of the learning-based approach. Despite the increased distributional complexity, NN maintains low MSE (below 2% for CL $\geq$ 85%) and narrow coverage ranges. By comparison, BB consistently overshoots the target coverage, while ILP suffers from high variance and instability, particularly with limited data.
The correctness rates again demonstrate the advantage of NN. At 90% CL, NN achieves nearly perfect correctness (99.8%), whereas ILP with 300 samples drops to 69.2%. Importantly, NN generates obstacle spaces that are consistently smaller than BB while maintaining high coverage reliability.
IV-D Sensor Uncertainty
For sensor uncertainty, we evaluate obstacles whose positions are estimated from noisy measurements. In this scenario, the true positions of two adjacent obstacles are observed through a sensor corrupted by Gaussian noise, resulting in uncertainty about their exact locations. The real obstacle distribution is obtained by sampling noisy measurements of the true positions and constructing the corresponding probabilistic obstacle space. The resulting distribution is illustrated in Fig. 8c.
[…]
Correctness rates remain above 98% for NN, ensuring that the generated obstacle spaces are reliable. Moreover, the obstacle space areas are significantly closer to the optimal solution compared with BB and ILP. This result keeps demonstrating NN’s ability to generalize well under noisy measurement conditions while avoiding excessive conservatism.
IV-E Overall Discussion
Across motion, behavior, and sensor uncertainty, the proposed NN consistently outperforms traditional BB and ILP methods. Three key observations can be drawn:
[…]
  • Reliability: NN maintains correctness rates above 93% in all scenarios, ensuring that the generated probabilistic obstacle spaces meet or exceed the desired confidence levels.
  • Efficiency: NN produces compact obstacle spaces that are significantly smaller than BB while avoiding the instability of ILP. This efficiency is particularly critical in path planning applications where over-approximation reduces safe space.
These results validate the effectiveness of the proposed learning-based framework in generating accurate and reliable probabilistic obstacle spaces under diverse forms of uncertainty.
V Generating speed analysis